A UV+VUV MULTIPHOTON IONIZATION SCHEME FOR OH RADICALS^{a}

JOSEPH M. BEAMES, \underline{FANG LIU} and MARSHA I. LESTER, Department of Chemistry, University of Pennsylvania, Philadelphia, PA 19104-6323.

OH radicals are of significant atmospheric interest as a dominant oxidizing agent in day-time tropospheric chemistry. In this study, a 1+1′ multiphoton ionization (MPI) scheme is employed to record rotationally-resolved spectra of OH radicals via the A^{2}Σ^{+} resonant intermediate state.^{a} UV excitation is used to prepare OH A^{2}Σ^{+}(v=1, J, F_{i}), which is subsequently ionized by a second photon of fixed frequency VUV (118.3 nm), generated by tripling the 355 nm output of a Nd:YAG laser. The mass-selected OH^{+} ion signal from 1+1′ MPI is detected using time-of flight mass spectrometry and compared with the laser-induced fluorescence (LIF) signal arising from OH A^{2}Σ^{+}-X^{2}Π(1,0) excitation. The MPI signal is observed over a range of UV+VUV total energies, corresponding to various OH A^{2}Σ^{+} (v=1, J, F_{i}) intermediate states, with relative intensities that differ considerably from LIF. The ion signal is enhanced relative to LIF at combined UV + VUV photon energies consistent with an autoionizing 3d Rydberg state that converges on the OH^{+} A^{3}Π ion core;^{b} direct ionization into OH^{+} X^{3}Σ^{−} is forbidden in a one-photon electric-dipole transition from OH A^{2}Σ^{+}. The MPI intensities have been quantified relative to LIF over the OH A^{2}Σ^{+}-X^{2}Π(1,0) region such that this scheme is now applicable for quantitative state-selective detection of OH X^{2}Π.

aThis research was supported by the Office of Basic Science of the Department of Energy.

aJ. M. Beames, F. Liu, M. I. Lester and C. Murray, J. Chem. Phys. \underline{134}, 241102 (2011).

bJ. D. Barr, A. De Fanis, J. M. Dyke, S. D. Gamblin, N. Hooper, A. Morris, S. Stranges, J. B. West, and T. G. Wright, J. Chem. Phys. \underline{110}, 345 (1999).